# Efficient Arbitrary Precision Acceleration for Large Language Models on GPU Tensor Cores

### v1

@@ Abstract.
Large language models (LLMs) have been widely applied but face challenges in efficient inference. While quantization methods reduce computational demands, ultra-low bit quantization with arbitrary precision is hindered by limited GPU Tensor Core support and inefficient memory management, leading to suboptimal acceleration. To address these challenges, we propose a comprehensive acceleration scheme for arbitrary precision LLMs. At its core, we introduce a novel bipolar-INT data format that facilitates parallel computing and supports symmetric quantization, effectively reducing data redundancy. Building on this, we implement an arbitrary precision matrix multiplication scheme that decomposes and recovers matrices at the bit level, enabling flexible precision while maximizing GPU Tensor Core utilization. Furthermore, we develop an efficient matrix preprocessing method that optimizes data layout for subsequent computations. Finally, we design a data recovery-oriented memory management system that strategically utilizes fast shared memory, significantly enhancing kernel execution speed and minimizing memory access latency. Experimental results demonstrate our approach’s effectiveness, with up to 13$\times$ speedup in matrix multiplication compared to NVIDIA’s CUTLASS. When integrated into LLMs, we achieve up to 6.7$\times$ inference acceleration. These improvements significantly enhance LLM inference efficiency, enabling broader and more responsive applications of LLMs.
@@ 1. Introduction
To address these challenges, we first propose a novel data format called bipolar-INT, which features a symmetric range and eliminates redundant sign bits, thereby reducing redundancy and facilitating parallel computing. Secondly, to overcome the limited precision support of GPU TCs, we achieve arbitrary precision MatMul through bit-level decomposition and recovery of matrices, saving memory while maintaining flexibility. Next, we introduce an efficient matrix preprocessing method that preprocesses input matrices by decomposing and reassembling them in advance, reducing communication costs and facilitating subsequent computations. Finally, we propose a memory management approach oriented by efficient data recovery, aiming to maximize the use of faster shared memory for tasks originally handled by global memory, greatly enhancing kernel execution speed. Ultimately, compared to NVIDIA’s MatMul acceleration design CUTLASS, our scheme achieves up to a 13$\times$ speedup in MatMul and up to a 6.7$\times$ inference acceleration when integrated into LLMs.
@@ 5.1. Arbitrary Precision Kernel Evaluation
In this subsection, we evaluate our arbitrary precision MatMul kernel design on both square and LLM-specific MatMul tasks, validating the effectiveness of our redundancy reduction and memory management techniques. We compare our designs, including W1A2 (1-bit weights, 2-bit activations), W2A2, and W3A4, with standard FP32 and FP16 MatMuls. We then benchmark against NVIDIA’s Tensor Cores-accelerated ultra-low bit MatMul (CUTLASS INT1 and CUTLASS INT4) to showcase our superior computational performance when using the same TCs. Finally, we compare with other TC-based MatMul acceleration techniques, such as APNN-TC (Feng et al., 2021), BSTC (Li et al., 2019), and BTC (Li and Su, 2020), to demonstrate the performance advantages of our work in low-bit arbitrary precision acceleration.
@@ 5.1.1. Square MatMul Performance
Compared to FP32 and FP16, our MatMul design shows increasing speedup as matrix size grows. For 4k×$\times$4k matrices, our W1A2 configuration achieves a remarkable 193$\times$ speedup over FP32 and about 70$\times$ over FP16. When compared to CUTLASS, which is limited to int1 and int4 precisions due to TC constraints, our design demonstrates superior performance. For 4k×$\times$4k matrices, our W1A2 configuration outperforms CUTLASS INT4 by more than 13$\times$. Notably, both our W1A2 and W2A2 configurations surpass CUTLASS INT1 in performance, despite not having a bit-width advantage. Our W1A2 configuration is 5.5$\times$ faster than CUTLASS INT1, while our W2A2 configuration is 3.5$\times$ faster. Even our W3A4 configuration, which has a significantly wider bit-width, approaches CUTLASS INT1 performance, achieving approximately 88% of its speed.
@@
While APNN-TC slightly outperforms our design for smaller matrices, our approach shows significant advantages for matrices 1k×$\times$1k and larger, with our W1A2 and W2A2 configurations achieving 44$\times$ and 50$\times$ speedups, respectively. This is particularly relevant for LLMs, where large matrix operations are more common. Furthermore, our design offers true arbitrary precision support, including configurations like W3A2 and W3A4, which APNN-TC cannot handle. This demonstrates not only superior performance but also greater flexibility, especially beneficial for the diverse computational needs of LLMs.
@@ 5.1.2. LLM-specific MatMul Performance
Our results show that while the acceleration effect slightly decreases for non-square matrices compared to square matrices, our approach still demonstrates outstanding performance. For instance, our W1A2 configuration achieves speedups of 91.2$\times$, 98.1$\times$, and 102$\times$ for the three matrix sizes tested, significantly outperforming both FP16 and CUTLASS implementations. Notably, our W1A2 and W2A2 configurations surpass CUTLASS INT1 despite its bit-width advantage, while our W3A4 configuration outperforms CUTLASS INT4. This superior performance can be attributed to our matrix-splitting strategy that minimizes data redundancy and our efficient memory management scheme.
@@
Fig. 6 illustrates the performance comparison across all MatMul sizes in Llama-7B, using TOPS as the metric. When compared to APNN-TC, our design shows similar performance for smaller matrices (e.g., 1k×$\times$1k×$\times$128). However, for larger matrices more common in LLMs (e.g., 1k×$\times$10.75k×$\times$4k), our approach achieves over 10$\times$ speedup compared to APNN-TC. This disparity is due to our design being optimized specifically for the large matrix computations prevalent in LLMs, whereas APNN-TC’s shared memory allocation and GPU thread scheduling are more suitable for smaller MatMul operations. In summary, our MatMul kernel design demonstrates significant performance advantages in LLM-specific scenarios, particularly for larger, non-square matrices that dominate LLM computations. This makes our approach particularly well-suited for accelerating LLM inference tasks.
@@ 6. CONCLUSIONS
In this paper, we present a comprehensive scheme to accelerate inference of ultra-low bit quantized large language models (LLMs) on GPU. We first introduce a data format called bipolar-INT, which is suitable for symmetric quantization and conducive to parallel computation. On top of bipolar-INT, we propose an INT MatMul method applicable to arbitrary precision and design an efficient memory management system. Our evaluations demonstrate significant performance gains in both MatMul operations and LLM inference. For MatMul, our implementation achieves a 5.45$\times$ speedup compared to Nvidia’s CUTLASS and up to 43$\times$ speedup over existing solutions. In LLM inference, our approach yields 3.9-6.7$\times$ speedup over FP16 models and 1.2-2$\times$ speedup compared to ultra-low bit quantized models using CUTLASS as the kernel.

### v2

@@ Abstract.
Large language models (LLMs) have been widely applied but face challenges in efficient inference. While quantization methods reduce computational demands, ultra-low bit quantization with arbitrary precision is hindered by limited GPU Tensor Core support and inefficient memory management, leading to suboptimal acceleration. To address these challenges, we propose a comprehensive acceleration scheme for arbitrary precision LLMs. At its core, we introduce a novel bipolar-INT data format that facilitates parallel computing and supports symmetric quantization, effectively reducing data redundancy. Building on this, we implement an arbitrary precision matrix multiplication scheme that decomposes and recovers matrices at the bit level, enabling flexible precision while maximizing GPU Tensor Core utilization. Furthermore, we develop an efficient matrix preprocessing method that optimizes data layout for subsequent computations. Finally, we design a data recovery-oriented memory management system that strategically utilizes fast shared memory, significantly enhancing kernel execution speed and minimizing memory access latency. Experimental results demonstrate our approach’s effectiveness, with up to 2.4$\times$ speedup in matrix multiplication compared to NVIDIA’s CUTLASS. When integrated into LLMs, we achieve up to 6.7$\times$ inference acceleration. These improvements significantly enhance LLM inference efficiency, enabling broader and more responsive applications of LLMs.
@@ 1. Introduction
To address these challenges, we first propose a novel data format called bipolar-INT, which features a symmetric range and eliminates redundant sign bits, thereby reducing redundancy and facilitating parallel computing. Secondly, to overcome the limited precision support of GPU TCs, we achieve arbitrary precision MatMul through bit-level decomposition and recovery of matrices, saving memory while maintaining flexibility. Next, we introduce an efficient matrix preprocessing method that preprocesses input matrices by decomposing and reassembling them in advance, reducing communication costs and facilitating subsequent computations. Finally, we propose a memory management approach oriented by efficient data recovery, aiming to maximize the use of faster shared memory for tasks originally handled by global memory, greatly enhancing kernel execution speed. Ultimately, compared to NVIDIA’s MatMul acceleration design CUTLASS, our scheme achieves up to a 2.4$\times$ speedup in MatMul and up to a 6.7$\times$ inference acceleration when integrated into LLMs.
@@ 5.1. Arbitrary Precision Kernel Evaluation
In this subsection, we evaluate our arbitrary precision MatMul kernel design on both square and LLM-specific MatMul tasks, validating the effectiveness of our redundancy reduction and memory management techniques. We compare our designs, including W1A2 (1-bit weights, 2-bit activations), and W2A2, with standard FP32 and FP16 MatMuls. We then benchmark against NVIDIA’s Tensor Cores-accelerated ultra-low bit MatMul (CUTLASS INT1 and CUTLASS INT4) to showcase our superior computational performance when using the same TCs. Finally, we compare with other TC-based MatMul acceleration techniques, such as APNN-TC (Feng et al., 2021), BSTC (Li et al., 2019), and BTC (Li and Su, 2020), to demonstrate the performance advantages of our work in low-bit arbitrary precision acceleration.
@@ 5.1.1. Square MatMul Performance
Compared to FP32 and FP16, our MatMul design shows increasing speedup as matrix size grows. For 4k×$\times$4k matrices, our W1A2 configuration achieves a remarkable 35.6$\times$ speedup over FP32 and about 13$\times$ over FP16. When compared to CUTLASS, which is limited to int1 and int4 precisions due to TC constraints, our design demonstrates superior performance. Notably, our W1A2 configurations surpass CUTLASS INT1 in performance, despite not having a bit-width advantage. For 4k×$\times$4k matrices, our W1A2 configuration outperforms CUTLASS INT4 by more than 2.4$\times$, while our W2A2 configuration is 1.24$\times$ faster. Even our W2A2 configuration, which has a significantly wider bit-width, approaches CUTLASS INT1 performance, achieving approximately 50% of its speed.
@@
While APNN-TC slightly outperforms our design for smaller matrices, our approach shows significant advantages for matrices 1k×$\times$1k and larger, with our W1A2 and W2A2 configurations achieving 7.9$\times$ and 7.2$\times$ speedups, respectively. This is particularly relevant for LLMs, where large matrix operations are more common. Furthermore, our design offers true arbitrary precision support, including configurations like W3A2 and W3A4, which APNN-TC cannot handle. This demonstrates not only superior performance but also greater flexibility, especially beneficial for the diverse computational needs of LLMs.
@@ 5.1.2. LLM-specific MatMul Performance
Our results show that while the acceleration effect slightly decreases for non-square matrices compared to square matrices, our approach still demonstrates outstanding performance. For instance, our W1A2 configuration achieves speedups of 61.8$\times$, 63.6$\times$, and 79.7$\times$ for the three matrix sizes tested, significantly outperforming both FP16 and CUTLASS implementations. Notably, our W1A2 and W2A2 configurations surpass CUTLASS INT1 most of the time despite its bit-width advantage. This superior performance can be attributed to our matrix-splitting strategy that minimizes data redundancy and our efficient memory management scheme.
@@
Fig. 6 illustrates the performance comparison across all MatMul sizes in Llama-7B, using TOPS as the metric. When compared to APNN-TC, our design shows similar performance for smaller matrices (e.g., 1k×$\times$1k×$\times$128). However, for larger matrices more common in LLMs (e.g., 1k×$\times$10.75k×$\times$4k), our approach achieves over 6.7$\times$ speedup compared to APNN-TC. This disparity is due to our design being optimized specifically for the large matrix computations prevalent in LLMs, whereas APNN-TC’s shared memory allocation and GPU thread scheduling are more suitable for smaller MatMul operations. In summary, our MatMul kernel design demonstrates significant performance advantages in LLM-specific scenarios, particularly for larger, non-square matrices that dominate LLM computations. This makes our approach particularly well-suited for accelerating LLM inference tasks.
@@ 6. CONCLUSIONS
In this paper, we present a comprehensive scheme to accelerate inference of ultra-low bit quantized large language models (LLMs) on GPU. We first introduce a data format called bipolar-INT, which is suitable for symmetric quantization and conducive to parallel computation. On top of bipolar-INT, we propose an INT MatMul method applicable to arbitrary precision and design an efficient memory management system. Our evaluations demonstrate significant performance gains in both MatMul operations and LLM inference. For MatMul, our implementation achieves a 2.4$\times$ speedup compared to Nvidia’s CUTLASS. In LLM inference, our approach yields 3.9-6.7$\times$ speedup over FP16 models and 1.2-2$\times$ speedup compared to ultra-low bit quantized models using CUTLASS as the kernel.